OpenStudy (anonymous):

I'm new to Trigonometry Could anyone help me expand "ln 8x"?

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle \log(a\times b)=\log(a)+\log(b) }\)

OpenStudy (solomonzelman):

Same thing would apply to a natural log: \(\large\color{slate}{ \displaystyle \ln(a\times b)=\ln(a)+\ln(b) }\)

OpenStudy (solomonzelman):

instead of a and b, you have 8 and x.

OpenStudy (anonymous):

Thank You so much! My teacher isn't very good lol

OpenStudy (solomonzelman):

Oh, tnx:) I am a calc II student, barely can do any math:) that is very encouraging....

OpenStudy (anonymous):

Literally, easiest explanation I've ever seen. Thanks again!

OpenStudy (solomonzelman):

Surely:) yw

OpenStudy (solomonzelman):

Also, \(\large\color{slate}{ \displaystyle \ln}\) is same thing as \(\large\color{slate}{ \displaystyle \log_e}\) (log w/ base e) for example: \(\large\color{slate}{ \displaystyle \ln(x)}\) is same thing as \(\large\color{slate}{ \displaystyle \log_e(x)}\) (just remember that) and when you have just a log, without an unspecified base, then the base is 10. in other words: \(\large\color{slate}{ \displaystyle \log(A)=\log_{10}A}\)

OpenStudy (solomonzelman):

there are other rules.. if I be around, you can always tag me or anyone else on here, by putting @ in front of the useranem. bye

OpenStudy (solomonzelman):

in front of the *username

OpenStudy (anonymous):

okay! thank you again!

OpenStudy (solomonzelman):

yw